# MaxMin-RLHF: Towards Equitable Alignment of Large Language Models with Diverse Human Preferences

### v1

@@ 4 Proposed Approach: Alignment with Diverse Preferences
9:     $\pi_{t+1}\leftarrow\text{PPO-update}(F_{{\phi_{u}^{*}}}(\pi_{t})-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi_{t}||\pi_{\text{ref}}\big{]})$
@@
4.1 Learning a Mixture Distribution for Diverse Preferences
To deal with the diversity in preferences, we first propose learning a mixture distribution for preferences rather than a single preference distribution, which is prevalent in the current RLHF pipelines (Christian, 2020; Ouyang et al., 2022b). The intuition behind learning a mixture of preference models is rooted in the principles of statistical learning theory (Lindsay, 1995; Seidel, 2011) which states that a combination of multiple distributions can approximate a wide range of complex, real-world distributions more effectively than a single model. This concept is akin to the property for mixture models, which suggests that a mixture model can approximate any number of diverse user sub-populations to an arbitrary level of accuracy, given a sufficient number of components and appropriate parameterization (Goodfellow et al., 2016). Thus, the preference distribution as shown in (3.1) can be expressed as
$p({\mathbf{z}}^{\prime})=\sum_{u=1}^{|\mathcal{U}|}p_{\phi_{u}}({\mathbf{z}}^{\prime})\cdot\eta(u),$ (13)
where the event $\mathbf{{\mathbf{z}}}^{\prime}:=(\mathbf{y}_{w}\succ\mathbf{y}_{l}\mid\mathbf{x})$, $y_{w},y_{l}$ are the chosen and rejected response by the human sub-population-group $\mathcal{H}_{u}$. With the definition in (13), the log-likelihood maximization objective to learn the sub-population preference distributions $\{p_{\phi_{u}}\}_{u=1}^{|\mathcal{U}|}$ parametrized by $\Phi=[\phi_{1},\phi_{2},\cdots\phi_{|\mathcal{U}|}]$ (under the Bradley Terry preference model in (1)) and the $|\mathcal{U}|$ groups/clusters can be written as
$\mathcal{L}(\Phi)$ $=\sum_{{\mathbf{z}}^{\prime}\in\mathcal{D}}\log\sum_{u=1}^{|\mathcal{U}|}p_{\phi_{u}}({\mathbf{z}}^{\prime})\cdot\eta(u),$ (14)
$=\sum_{{\mathbf{z}}^{\prime}\in\mathcal{D}}\log\sum_{u=1}^{|\mathcal{U}|}\frac{\exp\left(r_{\phi_{u}}(\mathbf{y}_{w},\mathbf{x})\right)}{\exp\left(r_{\phi_{u}}(\mathbf{y}_{w},\mathbf{x})\right)+\exp\left(r_{\phi_{u}}(\mathbf{y}_{l},\mathbf{x})\right)}\cdot\eta(u).$
We learn the parameter $\Phi^{*}$ by maximizing the above log-likelihood objective. It is important to emphasize that mixture models do not require prior knowledge of user group classifications, i.e., which user belongs to which group, and can learn the distribution of preferences in an unsupervised manner. This capability underscores the generality and flexibility of mixture distributions in capturing the diversity of user preferences without explicit user group information.
@@ 4.1 Learning a Mixture Distribution for Diverse Preferences
Remark. If we have access to individual human sub-population information, we can directly learn the reward parameter via maximizing likelihood per user group. But often, in practice, they are hardly available. To address this challenge, we consider an expectation-maximization algorithm to learn a mixture of reward models summarized in Algorithm 2 which learns the $r_{\phi_{u}}$’s and the $|\mathcal{U}|$ clusters. We summarize the EM algorithm for reward learning in Algorithm 2.
Figure 4: (a) This figure shows a Robotic navigation scenario where the objective of the agent is to distribute goods among the two groups denoted by green and orange boxes. In the above figure, (b) shows the trajectory when only green user group preferences (they want the agent to visit them) are considered to decide the agent path. (c) shows the trajectory when only green user group preferences are considered to decide the vehicle path. (d) shows the result for our proposed formulation (cf. Definition 4.2), where goal is to maximize the social utility, which ensures a robust solution to satisfy all user group preferences.
4.2 MaxMin RLHF
We begin by noting that while we circumvent the sub-optimality in reward learning (cf. Lemma 3.2) sub-population specific reward models ${\phi}^{*}_{u}$ for all subsets in $\mathcal{U}$ using Algorithm 2, this does not address the ultimate goal of language model alignment. Since the eventual objective is to develop a single policy $\pi^{*}$ that accommodates diverse user preferences while avoiding bias towards specific groups such as minorities. To achieve that, we take motivation from the Egalitarian rule in social choice theory (Sen, 2017), which states that society should focus on maximizing the minimum utility of all individuals. Hence, we write our proposed alignment objective, which maximizes the social utility as follows {defi}(Alignment with Diverse Human Preferences) From the mixture of preference distributions in (14), we get $\{r_{\phi_{u}^{*}}\}_{u=1}^{|\mathcal{U}|}$ where $\phi_{u}^{*}$ denotes reward model parameter for each human subpopulation in $\mathcal{U}$. We define the alignment objective in such diverse settings as
$\pi^{*}_{\text{MaxMin}}\in\arg\max_{\pi}\Big{(}\min_{u}\mathbb{E}\big{[}{r_{\phi_{u}^{*}}}(\mathbf{y},\mathbf{x})\big{]}\Big{)}-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (15)
where the expectation is with respect to $\{\mathbf{x}\sim\mathcal{P},\mathbf{y}\sim\pi(\cdot~{}|~{}\mathbf{x})\}$.
Proof of Concept: An illustration of MaxMin RLHF. To illustrate the social utility of the Max-Min objective within the context of RLHF, in Figure 4, we present a simplified robotic scenario. In this scenario, the robotics agent’s goal is to distribute goods among two user groups $u_{1},u_{2}$ located at green and orange boxes, respectively. Each user group would prefer that the agent visit them and distribute goods. Hence, we define the reward for each user group $u_{i}$ is defined as follows: the agent receives a reward $r_{u}(s,a)=C_{u}$ if it reaches the goal state for the user $u$ and $r_{u}(s,a)=0$, otherwise. We demonstrate that in this basic scenario, our proposed max-min algorithm ensures that the agent serves all the user groups irrespective of the difference in utility between the user groups $\Delta_{C}=|C_{1}-C_{2}|$. Mathematically, from the Definition 4.2, we can write
$\pi^{*}_{\text{MaxMin}}\in\mathop{\mathrm{argmax}}_{\pi}\min_{u}F_{u}(\pi)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (16)
[…]
$\pi^{*}_{\text{MaxMin}}\in$ $\mathop{\mathrm{argmax}}_{\pi}\min_{u}\sum_{s,a}r_{\phi_{u}}(s,a)\cdot\lambda^{\pi}(s,a)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$
$\in$ $\mathop{\mathrm{argmax}}_{\pi}{\min_{u}\mathcal{C}_{u}\lambda^{\pi}(s_{u},a_{u})}-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (17)
where $\lambda^{\pi}$ denote the occupancy measure corresponding to policy $\pi$. Let us focus on the minimization objective $\min_{u}\mathcal{C}_{u}\lambda^{\pi}(s_{u},a_{u})$, where $(s_{u},a_{u})$ signifies the goal state and action for a user $u$, and $\mathcal{C}_{u}$ is the corresponding reward. When the policy $\pi$ does not visit the goal for user $u$, the value of $\lambda^{\pi}(s_{u},a_{u})$ is zero, making the minimum always zero. Therefore, the maximum objective aims to learn a policy that maximizes the reward with respect to the users $u$ until $\lambda_{u}>0$ for all users. Even when the difference $\Delta_{c}=|\mathcal{C}_{1}-\mathcal{C}_{2}|$ is significantly large, highlighting a disparity between majority and minority user groups, the max-min approach effectively learns a policy that respects the needs of both users, demonstrating its social utility as shown in Figures 4.
MaxMin RLHF Algorithm. Next, we introduce the algorithm to solve the MaxMin RLHF proposed in Definition 4.2 in Algorithm 1. After learning the reward in Step 1 for individual user sub-population $r_{u}$ using Algorithm 2, we next compute the minimum utility user group as $u_{\text{min}}\leftarrow\arg\min_{u}F_{u}(\pi)$. Once the minimum utility user subpopulation group is determined, we fine-tune the language model policy for that group by taking gradient steps at each iteration $t$ as
$\pi_{{t+1}}=\pi_{t}+\eta\nabla_{\pi}(F_{u_{\text{min}}}(\pi_{t})-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi_{t}||\pi_{\text{ref}}\big{]}),$ (18)
where $F_{u_{\text{min}}}(\pi_{t}):=\mathbb{E}_{\mathbf{x}\sim\mathcal{P},\mathbf{y}\sim\pi_{t}(\cdot~{}|~{}\mathbf{x})}\big{[}{r_{u_{\text{min}}}}(\mathbf{y},\mathbf{x})\big{]}$ and $\eta$ denotes the step size. Note that for the implementation in practice, we leverage the PPO algorithm (Schulman et al., 2017) as commonly used in existing RLHF pipelines (Stiennon et al., 2022b; Christian, 2020; Ouyang et al., 2022b).
Algorithm 2 Learning Mixture Distribution with EM Algorithm
@@ 4.2 MaxMin RLHF
$\mathcal{H}_{u}:=\ arg\max_{u\in{1,\cdots,|\mathcal{U}|}}\prod_{(\mathbf{x},\mathbf{y_{1}},\mathbf{y_{2}},h)\in\mathcal{D}}w(\phi_{u},\mathbf{x},\mathbf{y_{1}},\mathbf{y_{2}})$
@@
4.3 Connection with DRO and General Utility RL
In this subsection, we highlight the generality of our proposed formulation beyond social utility design to broader formulations of general utility reinforcement learning (Zhang et al., 2020) and distributionally robust optimization (Duchi et al., 2018). We discuss then in details follows.
1. Distributionally Robust Optimization: We begin by noting that once we have learned the mixture of reward models $r_{{\phi}_{u}^{*}}$ for all user-groups $\mathcal{H}_{u}\in\mathcal{U}$ using Algorithm 2, we can formulate the problem of learning a robust policy as a distributionally robust optimization (DRO) (Ben-Tal et al., 2013; Duchi et al., 2018). In DRO, the objective is to maximize the worst-case expected utility over an uncertainty set of distributions. In our scenario, this boils down to constructing an uncertainty set over the distribution of rewards $r\in\mathcal{Q}$. Now, expanding the RL fine-tuning objective in the equation
$\pi^{*}_{\text{MaxMin}}\in\arg\max_{\pi}\min_{r\in\mathcal{Q}}\mathbb{E}_{r(y,x)}\rho_{\pi}(y,x)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (19)
where $\rho_{\pi}(y,x)=\mathcal{P}(x)\cdot\pi(y~{}|~{}\mathbf{x})$ as obtained from equation (3). The uncertainty set $\mathcal{Q}$ is typically obtained by constructing a distribution across the reward distributions $r$, however, this may lead to an overtly pessimistic solution. Hence, recent literature (Sagawa et al., 2020) has shown leveraging the prior knowledge of user-group information, can mitigate this by constructing the uncertainty set $\mathcal{Q^{\prime}}$ in terms of these groups $\mathcal{H}_{u}\in\mathcal{U}$, also referred to as the group distributionally robust optimization (Group-DRO). In Group DRO, the uncertainty set over reward distribution $\mathcal{Q}$ is constructed by a mixture of $u$ groups as $\mathcal{Q^{\prime}}=\sum_{u}\delta_{u}r_{u},\delta_{u}\in\Delta_{u}$, where $\Delta_{u}$ represents the $u-1$ dimensional probability simplex of preference distributions. However, since the inner problem is a linear program and the optimum is attained at the vertex, the resulting group DRO problem boils down to solving
$\pi^{*}_{\text{MaxMin}}$ $\in\arg\max_{\pi}\min_{u}\mathbb{E}_{r_{u}(y,x)}\rho_{\pi}(y,x)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$ (20)
$\in\arg\max_{\pi}\min_{u}F_{u}(\pi)-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]},$
where $r_{u}(y,x)$ represents the reward distribution of the user sub-population group $u$. It is important to note that the group-DRO formulation boils down to the Max-Min RLHF formulation under this selection of uncertainty set under the distributional reward setting. This highlights that our proposed MaxMin RLHF provides a socially fair and distributionally robust solution to alignment problem with diverse user groups.
2. General Utility Reinforcement Learning: Next, we highlight the connection of our proposed formulation to general utility reinforcement learning as proposed in (Zhang et al., 2020, 2021). Zhang et al. (2020) first proposed the notion of general utility reinforcement learning that goes beyond the standard setting of cumulative reward maximization to more general utilities where standard notions of Bellman’s equation, the value function, etc fail. Thus in our formulation with diverse user preferences, once we have learned the reward models $r_{\phi_{u}^{*}}$ for all user-groups $\mathcal{H}_{u}\in\mathcal{U}$ using Algorithm 2, we can formulate the general utility objective ${G}_{u}({\pi})$ as the minimizer over the value-functions of user groups $u$ as $G_{u_{\texttt{min}}}(\pi):=\min_{u}\langle\rho_{\pi},r_{u}\rangle-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi||\pi_{\text{ref}}\big{]}$. where, $\rho_{\pi}(y,x)=\mathcal{P}(x)\cdot\pi(y~{}|~{}\mathbf{x})$ as obtained from equation (3). With this, the general utility formulation boils down to
$\max_{\pi}\quad$ $G_{u_{\texttt{min}}}(\pi),$ (21)
Hence we can utilize the analysis developed in (Zhang et al., 2020, 2021) for the general utility RL to derive the theoretical convergence results for our proposed algorithm in this work.
@@ 5.1 Small Scale Experiments (with GPT-2): Sentiment and Conciseness Alignment
Figure 5: (Alignment with MaxMin RLHF). This figure shows the performance of our proposed MaxMin RLHF algorithm for the preference dataset described in Figure 2. The task is to align a language model to generate positive sentiment responses that are concise (of shorter token length) in nature. The top row shows the sentiment score (higher is better) distribution of the language model before alignment (left) and after alignment (right). The bottom row shows the conciseness score (lower is better) distribution of the language model before alignment (left) and after alignment (right). We note that MaxMin-RLHF aligned language model can generate highly positive sentiment sentences and satisfy the conciseness criteria. This shows alignment with both the majority and minority preferences.
@@ 5.2.1 Main Results
Impossibility of Single Reward Model. When the user groups are biased in-terms of the representation of the user-groups (divided into majority and minority groups based on the preference dataset), the single reward model fails to capture the preferences of minority user groups. We test on preference dataset P1A/P1B representing two user groups and adjust the ratio of the number of users from group P1A and group P1B. Table 1 summarizes the accuracy for the majority group and minority group, as well as the accuracy of the total data. Here, low evaluation accuracy is an indication that the eventual alignment with the minority user group will be poor after the PPO step since the reward model itself is not accurate.
@@
Reward Learning with EM (Algorithm 2). Following the procedures from Algorithm 2) in the experiment setup, we learn a mixture of reward models with EM algorithm on all three datasets until convergence, as shown in Figure 7. From the results in Figure 7, we note that after the fourth iteration, all users are clustered correctly, meaning the mixture preference model successfully converges we successfully learn diverse groups of users with diverse preferences.
MaxMin RLHF Alignment. We evaluate the performance of our MaxMin-RLHF alignment against conventional single-reward RLHF trained on preference datasets with diversity and representational bias. Our baselines include ratios of 1, 2, 6, and 10, the same setting as discussed in Table 1 for illustration of the representational bias in preference data. Following Jang et al. (2023), we use the same $50$ instances from Koala evaluation(Geng et al., 2023) and test the model’s ability to generate answers in different groups of users’ preferences. We run pairwise evaluations by GPT-4 using AlpacaFarm codebase(Dubois et al., 2023) and use the win rate to the base Tulu2-7B model as the metric. Our results in Table 2 and Table 3 show that MaxMin alignment keeps a high win-rate in comparison with conventional RLHF methods trained by PPO with a single reward model. Our MaxMin RLHF approach achieves an average improvement of more than 16% in win-rates over conventional RLHF algorithms and improves the win-rate (accuracy) for minority groups by over 33% without compromising the performance of majority groups. Table 2 illustrates that as the minority group’s representation in the preference data diminishes, the performance of conventional RLHF algorithms significantly deteriorates for them, whereas, our proposed Max-Min approach maintains a balance, respecting the performance of both groups by aligning with both.
@@
Method P1A P1B Average
MaxMin 57.50 60.00 58.75
1:1 56.00 51.97 53.99
2:1 57.78 44.00 50.89
6:1 54.81 48.00 51.41
10:1 55.11 45.08 50.10
Method P2A P2B Average
MaxMin 54.50 56.00 55.25
1:1 53.73 54.00 53.87
2:1 55.55 51.72 53.64
6:1 52.14 49.40 50.77
10:1 53.96 45.98 49.97
Table 3: Pairwise winrate (%) on P1-P2 using GPT-4. The table illustrates that as the minority group’s representation in the preference data diminishes, the performance of conventional RLHF algorithms significantly deteriorates for them, favoring alignment with the majority. In contrast, our proposed Max-Min approach maintains a balance, respecting the performance of both groups by aligning with both.
@@ 6 Conclusions
In this work, we critically examine the limitations of the conventional single-reward Reinforcement Learning from Human Feedback (RLHF) framework, particularly its insufficiency in addressing the diversity of human preferences, leading to an impossibility result for alignment with diverse preferences. To achieve a socially fair alignment in diverse human preference settings, we introduce a novel MaxMin-RLHF approach, which learns a max-min policy over mixture of reward functions to achieve a more equitable model alignment. Our experiments demonstrate the effectiveness of MaxMin-RLHF in producing socially fairer outcomes, highlighting the need for more inclusive strategies in RLHF methodologies.
@@ Appendix B A Detailed Context of Related Works
Reinforcement Learning from Human Feedback. RL methods, such as policy gradient, applied to train language models for long-form generation Cho et al. (2018). The current RLHF approaches (Stiennon et al., 2022b; Ziegler et al., 2020; Zhu et al., 2023) involve training a reward model based on human preference feedback and then fine-tuning the language model using proximal policy optimization (PPO) (Schulman et al., 2017). The PPO algorithm helps to learn a model that produces responses that maximize the reward (Ouyang et al., 2022b; Bai et al., 2022a). Besides PPO, DPO (Direct Preference Optimization, Rafailov et al. (2023)) directly trains the large language model using human preferences without training the reward model. A self-play-based approach such as SPIN (Chen et al., 2024) is similar to DPO but has an iterative framework. However, most of the existing alignment approaches only consider the average preference by human annotators and ignore the inherent diversity among human preferences (Casper et al., 2023; Kaufmann et al., 2023). A number of theoretical studies have analyzed the efficiency and benefits for reinforcement learning using preference data Ji et al. (2023b); Zhang et al. (2023); Li et al. (2023). Chakraborty et al. (2024) proposed a bilevel reinforcement learning framework for policy alignment. Recently (Santurkar et al., 2023) created a dataset for evaluating the alignment of language models with $60$ US demographic groups over a wide range of topics and found substantial misalignment between a selanguage models and those groups. It emphasizes the criticality of considering diversity while performing alignment.
@@
Diversity in Human Preferences. Here, we briefly review the literature highlighting the reasons for diversity in the context of LLMs. Diverse human preferences stem significantly from various factors related to social and cultural backgrounds (Aroyo et al., 2023b, a; Denton et al., 2021a). The key factors contributing to this diversity include (i) socio-demographic backgrounds, including race, ethnicity, age, and gender shape preferences. Gender differences, for example, influence sensitivity to online content, with women facing more online harassment Vogels (2021). (ii) Personal bias and context subjectivity, which affects the human preferences for controversial topics in interpreting language and divisive themes (Denton et al., 2021b; Sandri et al., 2023)). (iii) Imperfect preferences, which arises due to variations in expertise, training, or quality control leading to diverse preferences, with certain content inaccurately considered offensive by some groups (Sandri et al., 2023). (iii) Linguistic ambiguity & missing context, could lead to diversity because of words or phrases with multiple possible interpretations and without clear context (Sandri et al., 2023; Denton et al., 2021b; Sap et al., 2022). These factors collectively underscore the complexity of aligning LLM outputs with the diverse preferences of human users, demonstrating the importance of recognizing and addressing the multifaceted nature of user feedback.

### v2

@@ 4 MaxMin-RLHF: One Possibility
9:     $\pi_{i+1}\leftarrow\text{PPO-update}(F_{r_{\phi_{u}^{*}}}(\pi_{t})-\beta\mathbb{D}_{\textrm{KL}}\big{[}\pi_{t}||\pi_{\text{ref}}\big{]})$
@@
MaxMin RLHF. If we have access to individual human sub-population rewards, we can go directly to solve the optimization problem in (14) with the algorithm summarized in Algorithm 1. But often, in practice, they are hardly available. To address this challenge, we consider an expectation-maximization algorithm to learn a mixture of reward models summarized in Algorithm 2 which learns the $r_{\phi_{u}}$’s and the $|\mathcal{U}|$ clusters.
We summarize the EM algorithm for reward learning in Algorithm 2.
Algorithm 2 Learning Rewards with EM Algorithm
@@
$u=\ arg\max_{u\in{1,\cdots,|\mathcal{U}|}}\prod_{(\mathbf{x},\mathbf{y_{1}},\mathbf{y_{2}},h)\in\mathcal{D}}w(\phi_{u},\mathbf{x},\mathbf{y_{1}},\mathbf{y_{2}})$
@@ 5.1 Small Scale Experiments (with GPT-2): Sentiment and Conciseness Alignment
Figure 4: (Alignment with MaxMin RLHF). This figure shows the performance of our proposed MaxMin RLHF algorithm for the preference dataset described in Figure 2. The task is to align a language model to generate positive sentiment responses that are concise (of shorter token length) in nature. We note that MaxMin-RLHF aligned language model can generate highly positive sentiment sentences and satisfy the conciseness criteria. This shows alignment with both the majority and minority preferences.
@@ 5.2.1 Main Results
Impossibility of Single Reward Model. When the user groups are biased (divided into majority and minority groups based on the preference dataset), the single reward model fails to capture the preferences of minority user groups. We test on preference dataset P1A/P1B representing two user groups and adjust the ratio of the number of users from group P1A and group P1B. Table 1 summarizes the accuracy for the majority group and minority group, as well as the accuracy on the total data. Here, low accuracy means that the alignment with the minority user group will be poor after the PPO step since the reward model itself is not accurate.
@@
Reward Learning with EM (Algorithm 2). Following the procedures in the experiment setup, we get similar and good results on all three datasets, as shown in Figure 6. From the results in Figure 6, we note that after the fourth iteration, all users are clustered correctly, meaning the mixture preference model successfully converges we successfully learn diverse groups of users with diverse preferences.
MaxMin RLHF Alignment. We further test the performance of our MaxMin-RLHF alignment method and compare it with the single reward
@@
RLHF models trained on biased datasets. Our baselines include ratios of 1, 2, 6, and 10, the same setting as discussed for Table 1. Following Jang et al. (2023), we use the same $50$ instances from Koala evaluation(Geng et al., 2023) and test the model’s ability to generate answers in different groups of users’ preferences. We run pairwise evaluations by GPT-4 using AlpacaFarm codebase(Dubois et al., 2023) and use the win rate to the base Tulu2-7B model as the metric. Our results in Table 2 and Table 3 show that MaxMin alignment keeps a high win rate while the models trained by PPO with a single reward model on biased datasets will have a relatively poor performance on the minority data representing minority user groups.
Method P1A P1B
MaxMin 57.50 60.00
1:1 56.00 51.97
2:1 57.78 44.00
6:1 54.81 48.00
10:1 55.11 45.08
Method P2A P2B
MaxMin 54.50 56.00
1:1 53.73 54.00
2:1 55.55 51.72
6:1 52.14 49.40
10:1 53.96 45.98
Table 3: Pairwise winrate (%) on P1-P2 using GPT-4.
@@ 6 Conclusions
In this work, we critically examine the limitations of the single-reward RLHF framework, particularly its insufficiency in addressing the diversity of human preferences, leading to an impossibility result for alignment with diverse preferences. To achieve a socially fair alignment in diverse human preference settings, we introduce a novel approach called MaxMin-RLHF, which learns a max-min policy over a distribution of reward functions to achieve a more equitable model alignment. Our experiments demonstrate the effectiveness of MaxMin-RLHF in producing socially fairer outcomes, highlighting the need for more inclusive strategies in RLHF methodologies.
Impact Statement
The primary objective of this work is to highlight the limitations of the existing alignment techniques in representing diverse opinions and preferences. Our research is one of the first to formally highlight the above limitation with mathematical and empirical demonstrations. Finally, our research demonstrates a first step towards equitable alignment with diverse preferences. The findings of our work would encourage and foster further research in the domain of alignment under diversity, ensuring the current AI models is not biased towards specific minority groups.
Acknowledgements
Chakraborty and Huang are supported by DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) 80321, National Science Foundation NSF-IIS-2147276 FAI, DOD-ONR-Office of Naval Research under award number N00014-22-1-2335, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, DOD-DARPA-Defense Advanced Research Projects Agency Guaranteeing AI Robustness against Deception (GARD) HR00112020007, Adobe, Capital One and JP Morgan faculty fellowships. Manocha and Bedi are supported by Army Cooperative Agreement W911NF2120076. Mengdi Wang acknowledges the support by NSF IIS-2107304, NSF CPS-2312093, ONR 1006977 and Genmab. We also thank Rui Yang and Han Zhao for pointing out a bug in the proof of Lemma 1 in the previous version.
@@ Appendix B A detailed Context of Related Works
Reinforcement Learning from Human Feedback. RL methods, such as policy gradient, applied to train language models for long-form generation (Cho et al., 2018). The current RLHF approaches (Stiennon et al., 2022b; Ziegler et al., 2020; Zhu et al., 2023) involve training a reward model based on human preference feedback and then fine-tuning the language model using proximal policy optimization (PPO) (Schulman et al., 2017). The PPO algorithm helps to learn a model that produces responses that maximize the reward (Ouyang et al., 2022b; Bai et al., 2022a). Besides PPO, DPO (Direct Preference Optimization, Rafailov et al. (2023)) directly trains the large language model using human preferences without training the reward model. A self-play-based approach such as SPIN (Chen et al., 2024) is similar to DPO but has an iterative framework. However, most of the existing alignment approaches only consider the average preference by human annotators and ignore the inherent diversity among human preferences (Casper et al., 2023; Kaufmann et al., 2023). A number of theoretical studies have analyzed the efficiency and benefits for reinforcement learning using preference data (Ji et al., 2023b; Zhang et al., 2023; Li et al., 2023). (Chakraborty et al., 2024) proposed a bilevel reinforcement learning framework for policy alignment. Recently (Santurkar et al., 2023) created a dataset for evaluating the alignment of language models with $60$ US demographic groups over a wide range of topics and found substantial misalignment between a selanguage models and those groups. It emphasizes the criticality of considering diversity while performing alignment.
@@
Diversity in Human Preferences. Here, we briefly review the literature highlighting the reasons for diversity in the context of LLMs. Diverse human preferences stem significantly from various factors related to social and cultural backgrounds (Aroyo et al., 2023b, a; Denton et al., 2021a). The key factors contributing to this diversity include (i) socio-demographic backgrounds, including race, ethnicity, age, and gender shape preferences. Gender differences, for example, influence sensitivity to online content, with women facing more online harassment (Vogels, 2021). (ii) Personal bias and context subjectivity, which affects the human preferences for controversial topics in interpreting language and divisive themes (Denton et al., 2021b; Sandri et al., 2023)). (iii) Imperfect preferences, which arises due to variations in expertise, training, or quality control leading to diverse preferences, with certain content inaccurately considered offensive by some groups (Sandri et al., 2023). (iii) Linguistic ambiguity & missing context, could lead to diversity because of words or phrases with multiple possible interpretations and without clear context (Sandri et al., 2023; Denton et al., 2021b; Sap et al., 2022). These factors collectively underscore the complexity of aligning LLM outputs with the diverse preferences of human users, demonstrating the importance of recognizing and addressing the multifaceted nature of user feedback.